OpenStudy (anonymous):

If f (x)=x^2 -4 and g (x)=4x+3 find in simplest form a) f (g (x)) and b) g (f (x))

OpenStudy (anonymous):

a) f(g(x)=f(4x+3)=(4x+3)^(2) -4 b) g(f(x)=g(x^(2)-4)=4(x^(2)-4)+3